1. Let the money supply be $200 and the velocity of money be 2.5. Also let LRAS be Y=50.

a.Write a function representing aggregate demand and give an equilibrium price level in the long run.

b.Now imagine that due to “animal spirits” velocity falls to 2. Show the effect of this in the short run on AS/AD. (Show all the curves.)

c. If the central bank wants to keep the price level unchanged after the drop in velocity, what should they do (this requires a number)? Show this on a graph of AS/AD.
Answer #1

a) MV=PT

200*2.5=P50

P=500/50

=10

b) velocity falls to 2,

200*2=50p

P=400/50

P=8

to ASI ADU free level AS2 13 14 P ADZ o Y, Y AB, AS